a body starts from rest and travels distances of 120, 300 and 180m in successive equal time interval of 12s. during each interval the body is uniformly accelerated. i. calculate the velocity of the body at the end of each successive time interval, ii. sketch a velocity time graph for the motion


1
Expert's answer
2021-06-18T11:21:56-0400

The 120m-interval:


"v_1=\frac{2d_1}{t}=20\text{ m/s}."

The 300m-interval:


"v_2=v_1+a_2t=v_1+\frac{2d_2}{t^2}·t=v_1+\frac{2d_2}{t}=70\text{ m/s}."

The 180m-inetrval:


"v_3=v_2+\frac{2d_3}{t}=100\text{ m/s}."

The velocity-time graph for the motion: